I need help solving this equation. Thank you ahead of time for your help.
5/6(x-1) - 2/3(2x+1) = 3

Please note that 5/6 and 2/3 are fractions but not sure how to make the line straight.

Answer by stanbon(75887)

I need help solving this equation. Thank you ahead of time for your help.
(5/6)(x-1) - (2/3)(2x+1) = 3
-----
Multiply thru by 6 to get rid of the denominators:
5(x-1) - 4(2x+1) = 18
5x-5 - 8x-4 = 18
-3x - 9 = 18
x + 3 = -6
x = -9
================
Cheers,
Stan H.